OpenStudy (anonymous):

Help with exponential equation: Help with exponential equations: After solve the following exponential equation, x takes the form of 4^n x^(x^0.5) = 0.5^(1/2) What is teh value of n? Any idea , thanks.

OpenStudy (anonymous):

The equation is the following: 

OpenStudy (xanthe):

Sorry bt wheres the unknown n in the equation?

OpenStudy (anonymous):

It says, " x takes the form of 4^n " So we must solve for x and then for n.

OpenStudy (xanthe):

Sorry dude idk:(

OpenStudy (xanthe):

Bt is it a logarithm or calculas qn?

OpenStudy (anonymous):

I think we're not supposed to use logarithms, i tried that and was useless.

OpenStudy (gorv):

take log both sides x^0.5(logx)=(1/2)log0.5

geerky42 (geerky42):

I don't think it is possible to solve it algebraically. On calculator, graph \(y = x^{x^{0.5}}\) and \(y = \sqrt{0.5}\) and then find the value of x where they intersect. (there are two intersections) Hope this help.